OpenStudy (camerondoherty):

Out of the 150 students at a summer camp, 72 signed up for canoeing. There were 23 students who signed up for trekking, and 13 of those students also signed up for canoeing. Use a two-way table to organize the information and answer the following question: Approximately what percentage of students signed up for neither canoeing nor trekking? 45% 68% 28% 42%

OpenStudy (camerondoherty):

I think you nee to subract 13 from canoeing and both of them up

OpenStudy (camerondoherty):

then you get 82

OpenStudy (camerondoherty):

Subtract tht from 150

OpenStudy (camerondoherty):

and get 68

OpenStudy (camerondoherty):

then have it at 68

OpenStudy (camerondoherty):

but im not sureeee....

OpenStudy (taylorterminate):

I think ur correct. c:

OpenStudy (camerondoherty):

i need to be sure tho lol

OpenStudy (taylorterminate):

Wait i is calculating

OpenStudy (camerondoherty):

Do i multiply 0.68 by 150?

OpenStudy (shamil98):

72 + 23 =

OpenStudy (camerondoherty):

95

OpenStudy (shamil98):

95/150 * 100 = ?

OpenStudy (camerondoherty):

but 13 of those were already catgorized in canoeing

OpenStudy (camerondoherty):

so its 72+10

OpenStudy (shamil98):

thats irrelevant

OpenStudy (camerondoherty):

huh?

OpenStudy (shamil98):

95 students signed up for trekking or canoeing or both

OpenStudy (camerondoherty):

ohhhhhhh ok

OpenStudy (shamil98):

it says neither canoeing nor trekking, but it didnt specify if both so yes 72 + 10 82 150 - 82 = 68 68 out of 150 68/150 = 0.45 0.45 x 100 = 45 so 45%.

OpenStudy (shamil98):

make sense?

OpenStudy (camerondoherty):

Yes it does thank you c: and sry for da kinda late response lol my internet didnt show u replied

OpenStudy (taylorterminate):

lol

OpenStudy (shamil98):

alrighty.